OpenStudy (icalibear):

write an equation in point slope form of the line through point J(4, 1) with slope -4. My answer is y = -4x + 1 is that correct?

OpenStudy (anonymous):

No your equation is slope- intercept form . Point slope form is y – y1 = m(x – x1)

OpenStudy (icalibear):

So it's y - 1 = -4(x - 4)?

OpenStudy (anonymous):

yes sir that is correct

OpenStudy (icalibear):

thx